OpenStudy (anonymous):

Find the limit as x goes to infinity of x^20/e^x ?

OpenStudy (anonymous):

exponential grows faster than any polynomial

OpenStudy (xapproachesinfinity):

so if the denominator grow faster to the infinity what would the limit be?

OpenStudy (xapproachesinfinity):

you might want to recall that limit as x tends to inf of (x/e^x) is zero

OpenStudy (xapproachesinfinity):

if you did L'hospitals rule this would be much easier to take care of

OpenStudy (anonymous):

I don't understand why that is wouldn't that give your inf/inf

OpenStudy (anonymous):

really? you would have to do it like 20 times! easier to eyeball it

OpenStudy (anonymous):

Ohh I get it know thank you guys so much

ganeshie8 (ganeshie8):

Notice that "n" applications of lopital reduces the numerator of fraction to a constant, however not much happens in the denominator

ganeshie8 (ganeshie8):

\[\rm \lim\limits_{x\to \infty} \frac{x^{20}}{e^x} \stackrel{\color{red}{\text{LH 20 times}}}{\color{red}{\leadsto \leadsto\leadsto\leadsto}} \lim\limits_{x\to \infty} \frac{C}{e^x} \]

ganeshie8 (ganeshie8):

**nothing happens to the denominator, its no longer in indeterminate form, you can take the limit

ganeshie8 (ganeshie8):

exponential growth always overtakes ANY polynomial growth if you want long enough 

ganeshie8 (ganeshie8):

*if you wait long enough

OpenStudy (anonymous):

nice latex \[\rm \lim\limits_{x\to \infty} \frac{x^{20}}{e^x} \stackrel{\color{red}{\text{LH 20 times}}}{\color{red}{\leadsto \leadsto\leadsto\leadsto}} \lim\limits_{x\to \infty} \frac{C}{e^x}\]

OpenStudy (anonymous):

\[\rm \lim\limits_{x\to \infty} \frac{x^{20}}{e^x} \stackrel{\color{red}{\text{LH 20 times}}}{\color{red}{\leadsto \leadsto\leadsto\leadsto}} \lim\limits_{x\to \infty} \frac{20!}{e^x}\]